OpenStudy (anonymous):

Two numbers added together equal '6' One number divided by the other numbers equal '4' what are the numbers?

OpenStudy (anonymous):

xkeh...you not answering this?

OpenStudy (xkehaulanix):

Haha, I hit a bump. And got distracted. You go ahead.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You have two numbers and you don't know what they are. Let them by a and b. This is what we know about them:\[a+b=6\]and\[\frac{a}{b}=4\](it won't matter whether we divide a by b, or b by a). From the second equation, you have for a,\[a=4b\]which you can substitute into the first equation,\[a+b=(4b)+b=5b=6\]That is,\[5b=6 \rightarrow b=\frac{6}{5}\]Well, if b=6/5, the first equation says (you could use the second, but it would be messier)\[a+b=a+\frac{6}{5}=6 \rightarrow a=6-\frac{6}{5}=\frac{30-6}{5}=\frac{24}{5}\]

OpenStudy (anonymous):

So your numbers are\[a=\frac{24}{5},b=\frac{6}{5}\]which you can check.

OpenStudy (anonymous):

wow thankyou, im trying to understand and read it myself. give me time :)

OpenStudy (anonymous):

No worries. If you need help, ask anyone around. Maybe you'd like to become fans of xkehaulanix and myself? Click our 'Become a fan' buttons... ;)

OpenStudy (anonymous):

haha, iv'e already beat you to it ;)

OpenStudy (anonymous):

Awesome...cheers :)

OpenStudy (anonymous):

is there a way where you can ask people question without having everyone see? i wana ask more questions but don't want to seem dumb. hahah :)

OpenStudy (anonymous):

lol...dumb. Don't be silly. You can keep asking questions in a thread, like this - people will usually not bother to look if they're here to answer questions...people look for a red "no replies yet" thing in the left-hand box.

OpenStudy (anonymous):

By the way, no one knows who you are, so you don't have to worry about how you sound.

OpenStudy (anonymous):

Haha, well i guess you're right :) it's nice to finally fine something like this. i always look for a 'online math tutor' but this should do.

OpenStudy (anonymous):

Yeah, you'll be right here...and it's free...and a lot of us giving answers have degrees in mathematics.

OpenStudy (xkehaulanix):

Which is why you should look for the people who actually have degrees if you want to be sure you're getting right answers. It really is helpful. And you won't look dumb at all. It's more embarassing to answer a question completely wrong than to ask a question xD

OpenStudy (anonymous):

correct